Solve. 10x – 4 – 5x = 6 + 3x – 8
Answer by jai_kos(139)

Given 10x – 4 – 5x = 6 + 3x – 8
Get all the “x” terms to one side and the numbers to the other side
We get,
10x – 5x – 3x = 6 – 8 + 4
2x = 2
Divide the above equation by 2, we get
x = 1
Therefore the value of “x” is 1